OpenStudy (anonymous):

What is the value of 6!? Its a permutations test...I don't understand it... A:36 B:720 C:5040 D:46,656

OpenStudy (anonymous):

6

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

1 * 2 * 3 * 4 * 5 * 6

OpenStudy (yttrium):

Should also be n! = n (n-1)! 6! = 6(6-1)! = 6(5!)

OpenStudy (anonymous):

HELPPPPPPPP PLEASEEEEE

OpenStudy (anonymous):

1 * 2 * 3 * 4 * 5 * 6 = 720